OpenStudy (anonymous):

A quantity of energy equal to 8.90 × 10^12 joules was emitted from a nuclear reaction. What amount of mass was required to be converted to yield this amount of energy? (1 J = 1 kg times m squared over s squared ) 9.90 × 10^-5 kg 2.96 × 10^-4 kg 2.67 × 10^4 kg 8.01 × 10^5 kg

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

Use: \(\sf E=mc^2\)

OpenStudy (aaronq):

\((8.90 × 10^{12}~ kg~m^2/s^2)=m(3*10^8~m/s)^2\)

OpenStudy (anonymous):

Thanks!!